Notes from Random Hall House Meeting on 9/2/02

Random Hall MRNSA

September 4, 2002

1 MRNSA timeline

1. Random housemeeting (now!)

2. Other dorm meetings

3. DormCon meeting

4. Living group meeting (IFC, LGC, DormCon, Panhel)

5. Talk with UA and GSC

6. Talk with Tony Gray (RLA)

7. Write documents for admins.

8. Talk to larryben and Redwine

9. Talk to Corporation

10. Large public meeting

2 What Went Right

• Midway

• Dead week

• Good Orientation Leaders (fratboys especially)

• Always something to do (dorm activity)

• Got to meet people from all over campus

• Dinner w/profs

• Move-in and advisor on same day

• Moving frosh into temp rooms (get to meet them)

• Things going wrong ”builds community” (motivates people to change things)

3 What Went Wrong

• Not enough notice/organization of moving squads

• Didn’t get temp address info

• Need to get move in dates earlier

• Could only rank three housing choices

• Denise not sleeping

• Housemaster brunch — frosh too settled (maybe swap governments)

• Too many mailings

• Soft deadlines bad — have one comprehensive list

• Should have mailings online • Have students send comprehensive info mailing

• 2006 website hard to navigate

• Websites not up to date (use correction page)

• Need mailing on Athena (and give once here)

• Tell people about athena.dialup.mit.edu and SSH

• I3 CD not emphasized

• I3 CD needed Quicktime version many people didn’t have

• Quick poll – 14-0 used CD

– 12-2 watched videos

– 12.5-1.5 looked at other stuff

– 11-1 of those who watched videos liked Random’s

• Some dorms might need technical help with making videos

• Rush mailings should be emphasized

• Separate mailings by subject

• Have one point person for mailings

• Didn’t tell frosh RBA takes precedence and is binding

• RBA/language house role in the lottery

• Made RBA out to be more than it was

• Desks told frosh not to move

• Tech didn’t publish Daily Confusion every day

4 Other Suggestions

• Student run residence book

• Maybe have more info

• Make housing/roommate procedures clearer

• Make residence table more accurate

• Midway – Need dorms to take better advantage of it

– People should be better at answering questions

– Need more people

– Better intradorm communication

• Admissions material should mention dorm variety

• Make dorms less apathetic about dorm rush

• Maybe organize our own RBA

• Publicize how to get out of RBA

• Talk to GSC

• Help w/FSILG rush

• Guarantee housing for transfer students

• Have orientation run by office for student life

• Make parents go away – Better storage options?

– 40% of frosh in FPOPs — as easy for parents to come later • Grown-up agreement

• Maybe parents weekend isn’t necessary (get stats on how many attend)

• Write orientation schedule • Get stats

• Get important orientation dates

5 Issue Areas

5.1 Communication

• Administration to freshmen – Do better job of informing people of temp info

– Webpages should be well-organized and updated – Should have official online forum with guides

– Have upperclassmen contact them over the summer – Have fewer, better organized mailings

– Have a comprehensive deadlines list

– Have mailings online

– Send and have available info about Athena tips – Consider having students create residence book – Make housing/roommate procedures clearer

– Make table of residences info more accurate

– Admissions material should talk about dorm variety • administration to upperclassmen

– About what’s being told to freshmen

– Get move in dates for frosh earlier

– Should give important orientation dates

– Should give statistics

• About I3 — improve videos technically

• Publicize moving squads more

5.2 Timing

• Other orientation should not impinge on rush

• Housemaster brunch should be after frosh move

• Parents should leave, rather than making choices for frosh

5.3 General housing

• RACs should determine acceptable crowding

• Everyone should get early returns

• Should guarantee housing for transfer students

5.4 Summer housing lottery

• RBA took precedence and was binding

• Clarify RBA and language house role in lottery

5.5 Residence based advising

• Was advertised as being something other than it was • Should publicize how to get out of RBA

• Consider organizing own RBA

5.6 Orientation housing lottery

• Office should have extended hours after lottery

• RACs should check for mistakes before results announced • Should be able to rank as many dorms as wish

• Should be mandatory — must enter to squat dorm

• Results should come out earlier in the day

5.7 Making dorm rush work better

• Have a referral book with honest descriptions of all dorms • Have upperclassmen take frosh to other dorms

• Tech should publish Daily Confusion every day

• Desks shouldn’t tell frosh not to move

5.8 Orientation

• Perhaps should be run by Office of Student Life

• Get good Orientation Leaders

• Playfair was bad

• We should write a suggested Orientation schedule

5.9 Bad distractions during rush

• Internet access

• Frosh being drunk

• People they already know

5.10 Social engineering

• ”It would have been better if we scared our temp frosh.”

• Resentment from upperclassmen

• Help with FSILG rush